Factor completely. If the polynomial is prime, state this.
5p4q4 - 180

Answer by user_dude2008(1861)

^ means exponent
5p^4q^4 - 180 = 5(p^4q^4 - 36) = 5(p^2q^2-6)(p^2q^2+6)